OpenStudy (anonymous):

Mandy bought a bag of peanuts to share with her friends. • Trisha received 1/2 of the peanuts in the bag. • Vince received 1/4 of the peanuts Trisha received. • Ray received 1/3 of the peanuts Vince received. If Ray received 4 peanuts, how many peanuts were in the bag Mandy bought? A 13 B 32 C 96 D Not here

OpenStudy (anonymous):

c

OpenStudy (anonymous):

if Ray received 1/24 peanuts of what Trisha received. then 24*4 =96

OpenStudy (anonymous):

need explaining how i got it? or you were looking for an answer only?

OpenStudy (anonymous):

no, was wrong

OpenStudy (anonymous):

let me read more carefuul

OpenStudy (anonymous):

Ray received 1/3 of the peanuts Vince is mean vine 3 time of ray Vine = 4*3=12

OpenStudy (anonymous):

the answer in my key is C . Are you sure ?

OpenStudy (anonymous):

Trisha 1/2 than trisha twice time Vine Trisha =12*2=24

OpenStudy (anonymous):

isn't that what i got in the first place?

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

u r we com

OpenStudy (anonymous):

I think D

OpenStudy (anonymous):

trisha have 24 than the bag is 48 Mandy + 4+12+24=48 Mandy+40 = 48 Mandy =48 -40 Mandy = 8 The answer is D